OpenStudy (anonymous):

Simplify the radical expression. (radical)72 I don't understand how to do this. I found the square root, but it isn't an answer choice. Explain if you can! Thank you!

OpenStudy (akashdeepdeb):

\[\sqrt{72} = \sqrt{2 * 36} = \sqrt{2} * \sqrt{36} = 6 \sqrt{2}\]

OpenStudy (anonymous):

That's really helpful! Thank you! @AkashdeepDeb